OpenStudy (anonymous):

Here again who prove this without using mathematical induction method 1+2^2+3^2+4^2+⋯⋯+n^2=(n(n+1)(2n+1))/6

OpenStudy (mimi_x3):

without mathematical induction ? >_<

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

how is it possible ? >_<

OpenStudy (anonymous):

Use concept of limit.

OpenStudy (mimi_x3):

Oh sorry idk, probably someone else could help.

OpenStudy (jamesj):

To be clear, the ONLY way to prove this in general is with the Principle of Mathematical Induction. There is no proof of this identity using the concept of limits.

OpenStudy (anonymous):

Yes JamesJ it has.

OpenStudy (jamesj):

Show us.

OpenStudy (zarkon):

\[(n+1)^3-1=\sum_{i=1}^{n}[(i+1)^3-i^3]=\sum_{i=1}^{n}[i^3+3i^2+3i+1-i^3]\] \[=\sum_{i=1}^{n}[3i^2+3i+1]=\sum_{i=1}^{n}3i^2+\sum_{i=1}^{n}3i+\sum_{i=1}^{n}1\] \[=3\sum_{i=1}^{n}i^2+3\frac{n(n+1)}{2}+n\] Solve for \[\sum_{i=1}^{n}i^2\]

OpenStudy (jamesj):

Bhy is that first identity true: \[ (n+1)^3 - 1 = \sum ... \] and can you prove it without mathematical induction? Note also the proof of \( \sum_1^n i = n(n+1)/2 \) also required induction.

OpenStudy (zarkon):

the first identity is a telescoping sum for the sum of i s=1+2+...n s=n+n-1+...+1 2s=(n+1)+...+(n+1)=n(n+1) s=n(n+1)/2

OpenStudy (jamesj):

Ok, fair enough then.

OpenStudy (jamesj):

I stand corrected on the necessity of induction.

OpenStudy (jamesj):

But I still call bs on the notion of using limits. :-)

OpenStudy (amistre64):

cant win em all :) my goal is at least 5% so that I can stay statistically relevant lol